OpenStudy (anonymous):

simplify ^3√-343x^3 = my answer is -7x is this correct

OpenStudy (mathteacher1729):

I don't understand the question. Click the "equation" button and type your equation that way. :) It looks like it's asking... \[\sqrt[3]{-343x^3}\] but I want to make sure.

OpenStudy (anonymous):

yes it must be since 343 is a perfect cube. \[7^3=343\] so you have \[\sqrt[3]{-343x^3}=-7x\]